At time =0s the particle starts moving alopng x – axis. If its kinetic energy uniformly with time T , the net force acting on it must be prorportional to 
  1. T^1/2
  2. constant
  3. T
  4. 1/squareroot T

Grade:11

1 Answers

Nirmal Singh.
9 years ago
The Right question is this – – – -At time =0s the particle starts moving alopng x – axis. If its kinetic energy uniformly with time T , the net force acting on it must be prorportional to-----------

As per question we know that the K.E increase with time T

So K,E = k*T

0.5 m v2= k T

so v = sqrt (2k T / m)

a = dv / dt = ½ sqrt (2k.mT)

So Force F = ma

F = m *½ sqrt (2k.mT)

Hence Force is prportaional to the1/squareroot T

Regards,

Nirmal SIngh
Askiitians Faculty